# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Type Ic Broad-Lined Supernova ?

### v1

@@ 3.2 NIRES spectrum
The NIR spectrum of SN 2021lwz, obtained +12.6 days, rest frame, after peak is shown in Figure 4. It provides valuable information for comparison with several types of supernovae observed in the NIR. The spectrum is smooth and relatively featureless with an absorption feature centred at m (see gray vertical line with exclamation mark). If that feature was to be produced by He i, m, it’s velocity would be km s-1, i.e. about the same order as the ejecta velocity derived a few days before maximum light, and about two times larger than the photosphere velocities at late phase ( km s-1), as derived from the optical spectra analysis discussed in the following sections. At this late phase velocity, any He i, m absorption feature would be encountered at a wavelength position centred close to 1.0559 m as indicated by the black vertical line. For comparison, the high velocity km s-1, He i, m absorption feature possibly detected in the peculiar Type Ic SN 1998bw (Patat et al. 2001) is shown by the blue vertical line. Similar comparisons can be done with the He-Rich and He-Poor d template spectra of Stripped-Envelope SNe (SESNe) derived from a principal component analysis of Types IIb, Ib, Ic and Ic-BL by Shahbandeh et al. (2022), and with the days He-Rich spectrum of SLSN-I SN 2019hge (Yan et al. 2020). At higher wavelengths, these 3 spectra show spectral features which are not seen in the spectrum of SN 2021lwz, whose smooth features look more similar to those seen in the spectrum of SN 1998bw.
@@
To assess if SN 2021lwz could be a Type I SLSN, we also made comparisons with spectra (not shown here) of SN 2015bn at phases -9 days and +34 days (see Figure 11, Nicholl et al. 2016), of Gaia16apd at phase +0 days (see Figure 7, Yan et al. 2017) and of SN 2012il at phase (see Figure 9, Inserra et al. 2013). None of these spectra are at a phase similar to the NIR spectrum of SN 2021lwz, and the wavelength coverage is sometimes limited to make meaningful comparisons, but all in all, they are all looking different from the spectrum of SN 2021lwz.
@@ 3.3 P-Cygni velocity estimation
We use the Fritz data science platform to cycle through all of the available lines and find that Fe ii is the best one that matches the broad feature near 4515Å seen in some of the spectra at a redhsift z=0.065.
@@ 3.7 Bolometric light curve estimation
At the eight epochs where UVOT data is available we calculate the spectral luminosity (Lλ) for each UVOT filter measurement and GP interpolate filter value from the other filters. Short of the lowest filter value we create Lλ estimates based on the best fit blackbody to the UVOT-UV, UVOT-U and LT-u filter measurements. Long of the highest filter value we create Lλ estimates based on the best fit blackbody to the optical filter measurements. These extrapolation methods are also used in SuperBol software (Nicholl 2018). We then integrate over all wavelengths to create bolometric luminosity estimates, at each of the eight epochs, which form the bolometric light curve. These results are shown, modeled and discussed in section 4.2.
@@ 3.8 Photosphere geometry
The polarization degrees measured on the unpolarized star, BD52913, on the polarized calibration star, HD 127769, on the field star STAR 1 (shown in Figure 17), and on SN 2021lwz are all displayed in Table 8. The instrumental polarization () degree is of order of at each epoch. This component, which is retrieved in the parameter space of the Stokes and parameters, is removed from the components of the other sources ( correction). We use the polarized star HD 127769 to calculate the zero polarization angle () which is used to rotate the Stokes parameters from the ALFOSC FAPOL instrument reference frame to the sky reference frame in Equatorial coordinates. The polarization angles are counted positively from North to East. We use the polarimetry available on STAR 1, in the field of view of FAPOL (see Figure 17), to get an estimate of the polarized component of the Galaxy in the vicinity of SN 2021lwz (see values shown in bold in Table 8). These estimates are of the same order as their uncertainties suggesting a low level of the interstellar polarization () in this region of the sky. This result is consistent with the polarization degrees of the two stars HD 82885 () and HD 82635 () retrieved from the Heiles (2000) agglomeration catalog. These two stars are at distances lower than two degrees on the plane-of-sky from the central coordinate position of SN 2021lwz. The levels of polarization obtained on SN 2021lwz after correction using STAR 1 Stoke and parameters, are displayed in the last column of Table 8. The final values shown in bold are obtained after bias correction following the equation given in Wang et al. (1997):
@@ 4.1 Optical spectrum evolution and chemical composition
For SN 2021lwz, we have spectra available at 12 different epochs from -5.4 days to +29.1 days phase, rest-frame, relative to the moment of the maximum light. Table 1 presents the date of observation of the modeled high-resolution spectra together with their phases, and the and values of their best-fit models obtained using SYN++. The possibilities, constraints and uncertainties of the usage of SYN++ and its parameters are discussed in details in Könyves-Tóth (2022).
@@
Figure 10 shows the spectrum modeling of SN 2021lwz. In the top panel, the modeling of the -3.4d phase spectrum of SN 2021lwz can be seen. The best-fit and values can be read out from Table 1. The identified elements are C ii, O i, Na i, Mg ii, Si ii, P iii, Ca ii, Ti ii and Fe ii. Regarding this, SN 2021lwz is more similar to a normal Type Ic supernova rather than a SLSN. The pre-maximum spectrum of 2/3 of the SLSNe-I usually contains a W-shaped absorption blend between 4000 and 5000 Å, while the remaining 1/3 fraction of the SLSNe-I show similar features to SN 2015bn (see Könyves-Tóth 2022). However, the -3.4d phase spectrum of SN 2021lwz does not show resemblance to any of these SLSN-I subtypes.
@@
The bottom panel of Figure 10 shows the modeling of the +13.2d and the +29.1d phase spectra, which are quite similar to each other. By these phases, both the photospheric velocity and temperature decreased significantly ( and K), and besides Fe ii, Ca ii, Si ii, Mg ii, O i and C ii, He i was included in the best-fit model.
@@
In Table 9 in the Appendix, the best-fit local parameter values obtained in SYN++ are collected. The SYN++ modeling analysis suggests SN 2021lwz is a intermediate Type Ic/Ic-BL transient rather than a SLSN-I.
4.2 Bolometric Light curve modeling
@@
We compare semi-analytic models to the bolometric light curve to determine the most probable power source of the SN. The radioactive source and magnetar source models used by West et al. (2023, Sect. 7) are employed using the ejecta velocity estimate from Section 3.6.
By reparameterizing the 56Ni mass as a fraction of the ejecta mass we then force the fit to be physical. However, as shown by the dashed red line in Figure 11, the fit converges poorly.
4.3 Multiband Light curve modeling
@@
We perform inference on the multiband photometry using the Redback (Sarin et al. 2024) magnetar-driven supernova model including ejecta acceleration and non-vacuum dipole spin-down (Omand & Sarin 2024). We use the PyMultiNest sampler (Buchner et al. 2014) implemented in bilby (Ashton et al. 2019; Romero-Shaw et al. 2020), and sample in flux density with a Gaussian likelihood. The uncertainty presented is only the statistical uncertainty in the fits, and does not include systematic uncertainty inherent in the simplified one-zone Redback model.
@@
We sample the explosion time with a uniform prior of up to 100 days before the first observation, an extinction term with a uniform prior between 0 and 2, and a white noise term. We constrain the total rotational energy of the magnetar to be erg. We fixed the photospheric plateau temperature to a value of 8000 K to keep consistency with the value found in Section 4.1. The physical parameters and priors used are listed in Table 2.
@@
The fitted light curve is shown in Figure 12 and the posterior for the parameters in Figure 19 (see Appendix G). The model fits the peak of the light curve well in all bands but underestimates the flux in the UV bands from around 14 days post-peak. We find an initial luminosity erg s-1 and spin-down timescale one day, and an ejecta mass . These values are similar to that of the luminous fast-blue optical transient (LFBOT) AT 2020xnd (i.e. the ZTF source ZTF20acigmel; see e.g. Perley et al. 2021a), the ultra-stripped supernova (USSN) SN 2014ft (see Table 2 in Omand & Sarin 2024), and iPTF16asu (see Table 2; and Whitesides et al. 2017; Wang et al. 2019; Wang & Gan 2022), including a similar higher magnetar rotational energy of erg, but a lower explosion energy of erg. We find the braking index to be 4.5 and rule out vacuum dipole () with 95% confidence, similar to SN 2015bn and SN 2007ru (Omand & Sarin 2024).
@@
The distribution of ejecta velocities, peak photospheric temperatures, and several timescales is shown in Figure 13. The final ejecta scale velocity is around 11 000 km s-1, which is lower than the inferred photospheric velocity of 15 000 km s-1; this implies that most of the material is traveling at much lower velocity than the photosphere at early times, although by peak the photosphere has receded into the bulk of the material. The photospheric temperature at peak is around 24 000 K, slightly higher than the best-fit model temperature of 20 000 K inferred by the spectral analysis, although not significantly higher. The diffusion time of the supernova,
@@
which is 27 days post-explosion, or only 17 days post -band peak. This could imply that the final two observed spectra, at +26d and +29d, are taken during the early nebular phase, even though the spectrum shows no strong evolution between +8 and +29 days. This could also account for the discrepancy between the model and observed photometric points in the late-phase. However, it seems more likely that this simplistic estimate of the nebular timescale is not accurate in this circumstance, although it is not clear why this would be the case. Due to the large inferred gamma-ray opacity, the gamma-ray leakage timescale (Wang et al. 2015),
@@
is also very long, with most of the posterior being over 100 days. This means that non-thermal emission likely does not escape from the supernova over the timescale that we observed it. The high gamma-ray opacity here is inconsistent with a pulsar wind nebula (PWN) with low magnetization (Vurm & Metzger 2021), which would cause the non-thermal emission to be Comptonized and mostly emitted as hard x-rays and gamma-rays. The likely source of opacity for the PWN is photoelectric absorption, which dominates in the soft x-ray regime (see Figure 4 from Vurm & Metzger (2021)). This means that the PWN is likely synchrotron-dominated, and may be a strong radio source at longer timescales (e.g. Omand et al. 2018, 2025a).
@@
These parameters differ significantly from Gomez et al. (2022), and Gomez et al. (2024), who used the Modular Open Source Fitter for Transients code (MOSFiT; Guillochon et al. 2018) to fit several luminous supernovae and SLSNe, including SN 2021lwz, using a pure magnetar and magnetar nickel model, respectively. Gomez et al. (2022) infers a higher ejecta mass of 0.4 and a much lower photospheric velocity of 3000 km s-1. Gomez et al. (2024) infers much higher ejecta and 56Ni masses of 3 M and 0.8 M, respectively, while inferring a much less luminous pulsar of 1040 erg s-1, meaning the luminosity of the supernova is mostly powered by 56Ni in that scenario, in contrast to what we find in Section 4.2. While the values from Gomez et al. (2022) give a similar diffusion time to our inferred parameters, there is a strong tension between their derived photospheric velocity and the value we infer from the spectrum. They infer a lower value of erg s-1 for the initial pulsar luminosity but a spin-down timescale of 15 days. The difference in spin-down timescale, as well as their assumption of vacuum-dipole spin down, means the magnetar rotational energy is similar to our inferred parameters. Any discrepancies likely result from the assumption of constant velocity and decoupling of the magnetar spin-down and ejecta dynamics in the mosfit model, which has been shown via radiation hydrodynamics simulations to be unjustified for high-luminosity magnetars (Suzuki & Maeda 2021).
4.4 Host SED modeling
@@
The mass and the star-formation rate are comparable to common star-forming galaxies of that stellar mass (grey band in Figure 15; Elbaz et al. 2007)161616The relationship between star-formation and galaxy mass shown in Figure 15 is based on results from Brinchmann et al. (2004) and is valid at ., and they are also similar to those of the host galaxy populations of SNe from the PTF survey (Schulze et al. 2018) albeit at the very low-mass end of the observed distribution. The data relative to USSNe compiled in Table 10 are given in Appendix I.

### v2

@@ 3.2 NIRES spectrum
The NIR spectrum of SN 2021lwz, obtained +12.6 days, rest frame, after peak is shown in Figure 4. It provides valuable information for comparison with several types of supernovae observed in the NIR. The spectrum is smooth and relatively featureless with an absorption feature centred at m (see gray vertical line with exclamation mark). If that feature was to be produced by He i, m, it’s velocity would be km s-1, i.e. about the same order as the ejecta velocity derived a few days before maximum light, and about two times larger than the photosphere velocities at late phase ( km s-1), as derived from the optical spectra analysis discussed in the following sections. At this late phase velocity, any He i, m absorption feature would be encountered at a wavelength position centred close to 1.0559 m as indicated by the black vertical line. For comparison, the high velocity km s-1, He i, m absorption feature possibly detected in the peculiar Type Ic SN 1998bw (Patat et al. 2001) is shown by the blue vertical line. Similar comparisons can be done with the He-Rich and He-Poor template spectra at phase +13 days (d) of Stripped-Envelope SNe (SESNe) derived from a principal component analysis of Types IIb, Ib, Ic and Ic-BL by Shahbandeh et al. (2022), and with the days He-Rich spectrum of SLSN-I SN 2019hge (Yan et al. 2020). At higher wavelengths, these 3 spectra show spectral features which are not seen in the spectrum of SN 2021lwz, whose smooth features look more similar to those seen in the spectrum of SN 1998bw.
@@
To assess if SN 2021lwz could be a Type I SLSN, we also made comparisons with spectra of SN 2015bn at phases -9 days and +34 days (Nicholl et al. 2016b), of Gaia16apd at phase -2.5 days (Yan et al. 2017, and references therein) and of SN 2012il at phase (Inserra et al. 2013). None of these spectra are at a phase similar to the NIR spectrum of SN 2021lwz, and the wavelength coverage is sometimes limited to make meaningful comparisons, but all in all, they are all looking different from the spectrum of SN 2021lwz.
@@ 3.3 P-Cygni velocity estimation
We use the Fritz data science platform to cycle through all of the available lines and find that Fe ii is the best one that matches the broad feature near 4515 Å seen in some of the spectra at a redshift z=0.065.
@@ 3.7 Bolometric light curve estimation
At the eight epochs where UVOT data is available we calculate the spectral luminosity (Lλ) for each UVOT filter measurement and GP interpolate filter value from the other filters. Short of the lowest filter value we create Lλ estimates based on the best fit blackbody to the UVOT-UV, UVOT-U and LT-u filter measurements. Long of the highest filter value we create Lλ estimates based on the best fit blackbody to the optical filter measurements. These extrapolation methods are also used in SuperBol software (Nicholl 2018). We then integrate over all wavelengths to create bolometric luminosity estimates, at each of the eight epochs, which form the bolometric light curve. These results are shown, modelled and discussed in section 4.2.
@@ 3.8 Photosphere geometry
The polarization degrees measured on the unpolarised star, BD52913, on the polarized calibration star, HD 127769, on the field star STAR 1 (shown in Figure 18), and on SN 2021lwz are all displayed in Table 8. The instrumental polarization () degree is of order of at each epoch. This component, which is retrieved in the parameter space of the Stokes and parameters, is removed from the components of the other sources ( correction). We use the polarized star HD 127769 to calculate the zero polarization angle () which is used to rotate the Stokes parameters from the ALFOSC FAPOL instrument reference frame to the sky reference frame in Equatorial coordinates. The polarization angles are counted positively from North to East. We use the polarimetry available on STAR 1, in the field of view of FAPOL (see Figure 18), to get an estimate of the polarized component of the Galaxy in the vicinity of SN 2021lwz (see values shown in bold in Table 8). These estimates are of the same order as their uncertainties suggesting a low level of the interstellar polarization () in this region of the sky. This result is consistent with the polarization degrees of the two stars HD 82885 () and HD 82635 () retrieved from the Heiles (2000) agglomeration catalog. These two stars are at distances lower than two degrees on the plane-of-sky from the central coordinate position of SN 2021lwz. The levels of polarization obtained on SN 2021lwz after correction using STAR 1 Stoke and parameters, are displayed in the last column of Table 8. The final values shown in bold are obtained after bias correction following the equation given in Wang et al. (1997):
@@ 4.1 Optical spectrum evolution and chemical composition
For SN 2021lwz, we have spectra available at 12 different epochs from -5.4 days to +29.1 days phase, rest-frame, relative to the moment of the maximum light. Table 1 presents the date of observation of the modelled high-resolution spectra together with their phases, and the and values of their best-fit models obtained using SYN++. The possibilities, constraints and uncertainties of the usage of SYN++ and its parameters are discussed in details in Könyves-Tóth (2022).
@@
Figure 10 shows the spectrum modelling of SN 2021lwz. In the top panel, the modelling of the -3.4d phase spectrum of SN 2021lwz can be seen. The best-fit and values can be read out from Table 1. The identified elements are C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii. Regarding this, SN 2021lwz is more similar to a normal Type Ic supernova rather than a SLSN. The pre-maximum spectrum of 2/3 of the SLSNe-I usually contains a W-shaped absorption blend between 4000 and 5000 Å, while the remaining 1/3 fraction of the SLSNe-I show similar features to SN 2015bn (see Könyves-Tóth 2022). However, the -3.4d phase spectrum of SN 2021lwz does not show resemblance to any of these SLSN-I subtypes.
@@
The bottom panel of Figure 10 shows the modelling of the +13.2d and the +29.1d phase spectra, which are quite similar to each other. By these phases, both the photospheric velocity and temperature decreased significantly ( and K), and besides Fe ii, Ca ii, Si ii, Mg ii, O i and C ii, He i was included in the best-fit model.
@@
In Table 9 in the Appendix, the best-fit local parameter values obtained in SYN++ are collected. The SYN++ modelling analysis suggests SN 2021lwz is a intermediate Type Ic/Ic-BL transient rather than a SLSN-I.
To conclude, we point out that any difference between and can be explained with the presence of Non Local Thermodynamic Equilibrium (NLTE) effects. In the case of LTE, for each ion, however, in supernovae, NLTE effects usually play a role in the spectrum formation. Although SYN++ is not a NLTE code, like TARDIS (Kerzendorf and Sim 2014), which can appropriately take into account every difference from LTE, choosing a indicates that there is some divergence from the LTE. If we change the of a given ion in SYN++, the line ratios and strengths become different. We chose the best-fit model accordingly.
4.2 Bolometric light curve modelling
@@
We compare semi-analytic models to the bolometric light curve to determine the most probable power source of the SN. The radioactive source and magnetar source models used by West et al. (2023, Sect. 7) (Arnett 1982; Kasen and Bildsten 2010; Woosley 2010) are employed using the ejecta velocity estimate from Section 3.6. Our main conclusion is that the light curve can’t be fit with . The details of this analysis are given in Appendix G
There are no signs of narrow lines in the spectra indicating interaction, but shock breakout from interaction with extended material (Nakar and Sari 2010; Nakar and Piro 2014; Piro 2015) has been used to explain the light curve of the similar transient SN 2018gep (Ho et al. 2019; Pritchard et al. 2021). We do not explicitly use this model, but note that similar parameters (, 1052 erg, , and 1014 cm (Ho et al. 2019)) could likely also reproduce the light curve of SN 2021lwz.
4.3 Multi-band light curve modelling
@@
We perform inference on the multi-band photometry using the Redback (Sarin et al. 2024) magnetar-driven supernova model including ejecta acceleration and non-vacuum dipole spin-down (Omand and Sarin 2024). We use the PyMultiNest sampler (Buchner et al. 2014) implemented in bilby (Ashton et al. 2019; Romero-Shaw et al. 2020), and sample in flux density with a Gaussian likelihood. The uncertainty presented is only the statistical uncertainty in the fits, and does not include systematic uncertainty inherent in the simplified one-zone Redback model.
@@
We sample the explosion time with a uniform prior of up to 100 days before the first observation, an extinction term with a uniform prior between 0 and 2, and a white noise term. We constrain the total rotational energy of the magnetar to be erg. We fixed the photospheric plateau temperature to a value of 8000 K to keep consistency with the value found in Section 4.1. The physical parameters and priors used are listed in Table 2. Uniform priors are denoted by U and log-uniform priors are denoted by L.
@@
The fitted light curve is shown in Figure 11 and the posterior for the parameters in Figure 21 (see Appendix H). The model fits the peak of the light curve well in all bands but underestimates the flux in the UV bands from around 14 days post-peak. We find an initial luminosity erg s-1 and spin-down timescale one day, and an ejecta mass . These values are similar to that of the luminous fast-blue optical transient (LFBOT) AT 2020xnd (i.e. the ZTF source ZTF20acigmel; see e.g. Perley et al. 2021b), the ultra-stripped supernova (USSN) SN 2014ft (see Table 2 in Omand and Sarin 2024), and iPTF16asu (see Table 2; and Whitesides et al. 2017; Wang et al. 2019; Wang and Gan 2022), including a similar higher magnetar rotational energy of erg, but a lower explosion energy of erg. We find the braking index to be 4.5 and rule out vacuum dipole () with 95% confidence, similar to SN 2015bn and SN 2007ru (Omand and Sarin 2024).
@@
The distribution of ejecta velocities, peak photospheric temperatures, and several timescales is shown in Figure 12. The final ejecta scale velocity is around 11 000 km s-1, which is lower than the inferred photospheric velocity of 15 000 km s-1; this implies that most of the material is travelling at much lower velocity than the photosphere at early times, although by peak the photosphere has receded into the bulk of the material. The photospheric temperature at peak is around 24 000 K, slightly higher than the best-fit model temperature of 20 000 K inferred by the spectral analysis, although not significantly higher. The diffusion time of the supernova,
@@
which is 27 days post-explosion, or only 17 days post -band peak. This could imply that the final two observed spectra, at +26d and +29d, are taken during the early nebular phase, even though the spectrum shows no strong evolution between +8 and +29 days. This could also account for the discrepancy between the model and observed photometric points in the late-phase. However, it seems more likely that this simplistic estimate of the nebular timescale is not accurate in this circumstance, which could be because of small scale asymmetries or the density profile of the ejecta. Since Equations 49 are all derived using a uniform density profile, they could all underestimate the different timescales if the ejecta has a steeper profile (e.g. Suzuki and Maeda 2017). Due to the large inferred gamma-ray opacity, the gamma-ray leakage timescale (Wang et al. 2015),
@@
is also very long, with a lower limit of 50 days. This means that a significant amount of non-thermal emission likely does not escape from the supernova over the timescale that we observed it. The inferred lower limit on the gamma-ray opacity is 0.2 cm2 g-1, which is higher than the mean value of 0.06 cm2 g-1 for SLSNe-I (Gomez et al. 2024), although not a significant outlier. The high gamma-ray opacity here may be inconsistent with a pulsar wind nebula (PWN) with low magnetization (Vurm and Metzger 2021), which would cause the non-thermal emission to be Comptonized and mostly emitted as hard x-rays and gamma-rays. The likely source of opacity for the PWN is photoelectric absorption, which dominates in the soft x-ray regime (see Figure 4 from Vurm and Metzger (2021)). This means that the PWN is likely synchrotron-dominated, and may be a strong radio source at longer timescales (e.g. Omand et al. 2018, 2025a).
@@
These parameters differ significantly from Gomez et al. (2022), and Gomez et al. (2024), who used the Modular Open Source Fitter for Transients code (MOSFiT; Guillochon et al. 2018) to fit several luminous supernovae and SLSNe, including SN 2021lwz, using a pure magnetar and magnetar nickel model, respectively. Gomez et al. (2022) infers a higher ejecta mass of 0.4 and a much lower photospheric velocity of 3000 km s-1. Gomez et al. (2024) infers much higher ejecta and 56Ni masses of 3 M and 0.8 M, respectively, while inferring a much less luminous pulsar of 1040 erg s-1, meaning the luminosity of the supernova is mostly powered by 56Ni in that scenario, in contrast to what we find in Section 4.2. While the values from Gomez et al. (2022) give a similar diffusion time to our inferred parameters, there is a strong tension between their derived photospheric velocity and the value we infer from the spectrum. They infer a lower value of erg s-1 for the initial pulsar luminosity but a spin-down timescale of 15 days. The difference in spin-down timescale, as well as their assumption of vacuum-dipole spin down, means the magnetar rotational energy is similar to our inferred parameters. Any discrepancies likely result from the assumption of constant velocity and decoupling of the magnetar spin-down and ejecta dynamics in the mosfit model, which has been shown via radiation hydrodynamics simulations to be unjustified for high-luminosity magnetars (Suzuki and Maeda 2021).
4.4 Host SED modelling
@@
The mass and the star-formation rate are comparable to common star-forming galaxies of that stellar mass (grey band in Figure 14; Elbaz et al. 2007)161616The relationship between star-formation and galaxy mass shown in Figure 14 is based on results from Brinchmann et al. (2004) and is valid at ., and they are also similar to those of the host galaxy populations of SNe from the PTF survey (Schulze et al. 2018) albeit at the very low-mass end of the observed distribution. The data relative to USSNe compiled in Table 10 are given in Appendix J.
As a consistency check, we computed a plot similar to the one shown in Figure 13, without taking into account the emission lines. We find that the galaxy mass is almost unchanged. On the other hand the SFR departs from a factor 10 from the one discussed in the text and shown in Figure 14. The new reduced chi-square is also higher. All these points make us confident that, despite the low number of data points, the fit shown in Figure 13 including the emission-lines is better suited than a fit that would not include these lines.